Does there exist a plane targent to x
2 −2y
2 +2z
2 = 8 and which passes through
2x+3y+2z = 8, x−y+2z = 5? Justify your answer.
1
Expert's answer
2020-03-16T13:53:51-0400

"x^2-2y^2+2z^2=8"

The plane which passes through 

"2x+3y+2z=8\\
x-y+2z=5"

Point

"y=0\\
\left\{\begin{matrix}
 2x+2z=8\\
 x+2z=5
\end{matrix}\right.\\
x=3\\
z=1\\
A(3,0,1)"

directional vector line

"\vec{a}=\vec{n_1} \times\vec{n_2}\\
\vec{n_1}=(2,3,2)\\
\vec{n_2}=(1,-1,2)\\
\vec{a}=(\begin{vmatrix}
 3 & 2 \\
 -1 & 2
\end{vmatrix}; -\begin{vmatrix}
 2 & 2 \\
 1 & 2
\end{vmatrix};\begin{vmatrix}
 2& 3 \\
 1 & -1
\end{vmatrix})=\\
=(8;-2;-5)"

Take any point on

"x^2-2y^2+2z^2=8\\
M(x_0;y_0,z_0)\\
x_0^2-2y_0^2+2z_0^2=8\\
x_0^2=2y_0^2-2z_0^2+8"

The equation of the plane passing through the points "M, A" and the vector "\vec{a}"

"\begin{vmatrix}
 x-3 & y-0 &z-1\\
 x_0-3 & y_0-0&z_0-1\\
8&-2&-5
\end{vmatrix}=0\\
(x-3)\begin{vmatrix}
 y_0 & z_0-1\\
 -2 & -5
\end{vmatrix}-y\begin{vmatrix}
 x_0-3 & z_0-1 \\
 8 & -5
\end{vmatrix}+\\+(z-1)\begin{vmatrix}
 x_0-3 & y_0 \\
 8 & -2
\end{vmatrix}=0\\
(x-3)(-5y_0+2z_0-2)-y(-5x_0-8z_0+23)+\\
+(z-1)(-2x_0-8y_0+6)=0"

Answer: there plane a exist and have equation

"(x-3)(-5y_0+2z_0-2)-y(-5x_0-8z_0+23)+\\
+(z-1)(-2x_0-8y_0+6)=0"

where "x_0^2=2y_0^2-2z_0^2+8"